OpenStudy (toxicsugar22):

Find the indicated term in the product. (3x+x^3-7x^5)(1+4x-3x^2) for x^3

OpenStudy (toxicsugar22):

can u help me

OpenStudy (toxicsugar22):

can someone help

OpenStudy (toxicsugar22):

blexting can u help me

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

There is a long way and a short way. the long way is to multiply this out.. you would distribute 3 times.

OpenStudy (anonymous):

Or. you could look at how can I get x^3 by multiplying terms. If you are not that good at math we should do it the long way.

OpenStudy (anonymous):

(3x + x^3 - 7x^5) (1) = 3x + x^3 - 7x^5 (3x + x^3 - 7x^5)(4x) = 12 x^2 + 4x^4 - 28x^6 (3x + x^3 - 7x^5)(-3x^2) = -9x^3 - 3x^5 + 21x^7 Now we only look at the x^3 terms. I see a x^3 and -9x^3 so if you add those we get -8x^3

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

that is the answer

OpenStudy (anonymous):

Yes

OpenStudy (toxicsugar22):

thanks and can u help me with one more

OpenStudy (anonymous):

Sure

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

it bassicly the same thing

OpenStudy (toxicsugar22):

is that okay

OpenStudy (anonymous):

Yes

OpenStudy (toxicsugar22):

Find the indicated term in each product (4+2x-x^2)(2-3x+2x^2) for x^2

OpenStudy (anonymous):

(4 + 2x - x^2)(2) = 8 + 4x - 2x^2 (4 + 2x - x^2)(-3x) = -12x - 6x^2 + 3x^3 (4 + 2x - x^2)(2x^2) = 8x^2 + 4x^3 - 2x^4 do you understand how I get these?

OpenStudy (toxicsugar22):

yes

OpenStudy (anonymous):

OK.. now look at all the terms on the right. typically we would add all of these up to get the answer, but they only want the x^2 term

OpenStudy (anonymous):

So I see... -2x^2 and -6x^2 and 8x^2

OpenStudy (anonymous):

So add those up. -2x^2 - 6x^2 + 8x^2 = 0x^2 so actually you have no x^2 terms in the answer. So I guess I would say 0x^2

OpenStudy (toxicsugar22):

wow thanks i got it

OpenStudy (toxicsugar22):

i gave u a medal

OpenStudy (anonymous):

Thanks... good luck